OpenStudy (shamil98):

*Calculus* Car B is 30 miles directly east of Car A and begins moving west at 90 mph. At the same moment car A begins moving north at 60 mph. What will be the minimum distance between the cars and at what time t does the minimum distance occur ?

OpenStudy (psymon):

Grr, hate these kind of problems, haha. Okay, related rates x_X

OpenStudy (timaashorty):

The answer is B. the minimum distances.

OpenStudy (shamil98):

There are no answer choices, this is free response, please refrain from trolling my questions, tima :)

OpenStudy (timaashorty):

@awkwardpanda I LOVE YOU <3 LMAO MADE LAUGH SO HARD!

OpenStudy (timaashorty):

*me

OpenStudy (timaashorty):

Shamil I'm sorry, my duty is to help others in here if you don't appreciate my help, then others will.

OpenStudy (psymon):

Im sure you know @zepdrix lol

zepdrix (zepdrix):

thinkinggg :c

OpenStudy (psymon):

I dont know what to do with y, lol. I can get the set-up but with the stranded y variable there I dont know xD

OpenStudy (awkwardpanda):

It's 69.

zepdrix (zepdrix):

Still trying to figure this out, but side note: Why is car B going 90 miles per hour? that seems rather dangerous +_+

OpenStudy (awkwardpanda):

Best response ever.

OpenStudy (shamil98):

I don't know man, because simon said to D:

OpenStudy (psymon):

\[x^{2} + y^{2} = d^{2}\] \[2x \frac{ dx }{ dt }+ 2y\frac{dy}{dt}= 2s\frac{ds}{dt}\] x = 30, dx/dt = -90, dy/dt = 60 \[2(30)(-90) + 2y(60) = 2s\frac{ds}{dt}\] And thats all I got, haha.

OpenStudy (shamil98):

I was thinking of approaching it like so variable x is the distance car A travels in t hours, and variable y is the distance car B travels in t hours. and L is e the distance between cars A and B after t hours..

Directrix (directrix):

@Psymon . Don't spam! Common forms of spam include: •Posting a reply that is not relevant to the topic of discussion within that question. http://openstudy.com/code-of-conduct

OpenStudy (psymon):

How is that not relevant? :/

OpenStudy (luigi0210):

@Directrix I don't think that's spam..

OpenStudy (psymon):

My post was more relevant than much of theothercrap above -_-

OpenStudy (shamil98):



Directrix (directrix):

@awkwardpanda Don't be offensive, inappropriate, or creepy! So avoid: Sexual innuendo http://openstudy.com/code-of-conduct

OpenStudy (awkwardpanda):

-sigh- sorry.. -_- ;-;

OpenStudy (shamil98):

and to find the distance of L \[L = \sqrt{x^2 +(30-y)^2}\] and then because the travel is at a constant using this (distance traveled) = (rate of travel) (time elapsed)

OpenStudy (psymon):

I still dont see how I did any spam, im actually posting an idea to the problem, whether its anywhere near the correct answer or not. Otherwise you might as well report this entire question and everyone in it -_-

OpenStudy (psymon):

And how is that 30-y?

zepdrix (zepdrix):

That drawing is so confusing :( why is the `vertical component` x?

OpenStudy (shamil98):

and 60mph is car a so it would be x = 60t and car B went 90mph is y = 90t

OpenStudy (shamil98):

I've worked out this much so far..

OpenStudy (anonymous):

Start out with \[ z^2=x^2+y^2 \]So \[ 2zz'=2xx'+2yy' \]

OpenStudy (psymon):

Okay, at least I got that far, haha.

OpenStudy (shamil98):

Then substitutuing those values for the equation of L, right?..

OpenStudy (shamil98):

\[L =\sqrt{(60t)^2 +(30-90y)^2}\] \[=\sqrt{3600t^2 + (30-90y)^2}\]

OpenStudy (anonymous):

So you want\[ z'=0=\frac{xx'+yy'}{z}=\frac{90x+60y}{\sqrt{x^2+y^2}} \]

OpenStudy (anonymous):

Or\[ z' = 90x+60y \]

OpenStudy (shamil98):

Wouldn't i just differentiate with the chain rule?..

OpenStudy (anonymous):

can one of you guys help me when you're done helping him? http://openstudy.com/study#/updates/5258f3d5e4b002bdb08f5d62

OpenStudy (awkwardpanda):

Tima got dis.

OpenStudy (timaashorty):

Wio and Pysmon, That is so sweet, and kind of you two to take your time and help this kid. i hope sham apreciates your help at least.

OpenStudy (timaashorty):

Shamil you still confused ? need more help ?

OpenStudy (psymon):

Except I cant help. I only have an idea xD I can get stuff like wio did, but I dont know what to do with it.

OpenStudy (timaashorty):

@iambatman

OpenStudy (anonymous):

hollaaa

OpenStudy (shamil98):

I'll continue then ._. \[L' =(1/2)(3600t^2+(30-90t)^2)^-1/2 \]

OpenStudy (timaashorty):

You still tried Pysmon which is reeally great of you <3

OpenStudy (psymon):

But yes, I thought sham was somewhat on the right track. \[d = \sqrt{(30-90t)^{2}+ (60t)^{2}}\] we want ds/dt to be a minimum, so we differentiate and set the derivative = 0 to try and find critical points.

OpenStudy (psymon):

\[\frac{ ds }{ dt }= \frac{ 1 }{ 2 }((30-90t)^{2}+ (60t)^{2})^{-1/2}((2)(30-90t)(-90)+120t)\]

OpenStudy (shamil98):

\[(7200t^2 + 2(30-90t)(-90))\]

OpenStudy (awkwardpanda):

@TakeA.I.M.

OpenStudy (anonymous):

here is what i get.. L = [ (30-90t)^2 + (60t)^2 ] ^(1/2) the derivative with respect to time is: (30 (13t-3) ) / sqrt( 13t^2 - 6t + 1) (30 (13t-3) ) / sqrt( 13t^2 - 6t + 1) = 0, solve for t t = 0.23 hr i'll hand check the times 0.24 and 0.22 to make sure they merge at 0.23 :)

OpenStudy (shamil98):

\[= \frac{ 23,400t - 5400}{ 2\sqrt{3600t +(30-90t)^2} }\]

OpenStudy (shamil98):

= 0 if a/b= 0 then 23,400t - 5400 =0 23,400t = 5400 t \[\approx 23\]

OpenStudy (shamil98):

t = 0.23 = 13.8 mins x = 13.85 mi, then y = 20.77 mi so the shortest possible distance would be \[L \approx24.96 mi\]

OpenStudy (shamil98):

I just need someone to check my work now ._.

OpenStudy (anonymous):

yeah its right

OpenStudy (anonymous):

made me erase all my work ty takeaim

OpenStudy (psymon):

Yeah, id have never gotten it.

OpenStudy (shamil98):

Thanks, everyone, excluding the irrelevant posters tima and panda.. :) My brain is exhausted a bit now though aha xD again Thank you everyone ! :)

OpenStudy (awkwardpanda):

Glad to help :D @timaashorty

OpenStudy (timaashorty):

You are very welcome Shamil98 & Panda, I believe you did a brilliant job in helping him, and correcting my answer. Thank you.

OpenStudy (awkwardpanda):

Thank you, Tima :D now time for cake.

OpenStudy (timaashorty):

Let's go eat that cake hun (:

OpenStudy (shamil98):

wasn't this around the time where I got you two banned? lel

OpenStudy (timaashorty):

oh my, READING THIS MAKES ME LOL!

OpenStudy (timaashorty):

Yes it is !WOW 23 days ? ;o

OpenStudy (shamil98):

ikr o.o

OpenStudy (timaashorty):

Why can't you get chat ?

OpenStudy (awkwardpanda):

I'm on my iPod xD

OpenStudy (timaashorty):

awwh ;c